What is the pH of a solution made by dissolving 6.86 grams of calcium fluoride in enough water to make 3.0×102 mL of solution? The Ka for HF is 6.8x10–4.
Answer #1

Concentration of CaF2,c= (mass/molarmass)*(1000/volume in mL)

=(6.86 g/78(g/mole))*(1000/3.0*10^2 mL)

= 0.293 M

Let a be the dissociation of the weak acid
HF <---> H + + F-
initial conc. c 0 0
change -ca +ca +ca
Equb. conc. c(1-a) ca ca

Dissociation constant , Ka = ca x ca / ( c(1-a)
= c a^2 / (1-a)
In the case of weak acids a is very small so 1-a is taken as 1
So Ka = ca^2
==> a = √ ( Ka / c )
Given Ka = 6.8x10^-4
c = concentration = 0.293 M
Plug the values we get a = 0.0482

[H+]=ca= 0.293*0.0482=0.0141 M

pH= -log[H+]=-log0.0141=1.85